# Rui Lia,  Anyao Wangb   Mingqing Zhaic

### v2

@@ 4 Extremal graphs
Now, we consider the case where $g_{0}$ does not satisfy the inequality $43t_{0}\leq g_{0}<8t_{0}-$ 4. By a similar argument as in Claim 4.3, if $\mathcal{N}_{1}$ contains a non-tree component $H$ that is not a $K_{3}$, then we can transform $H\cup aK_{1,\lfloor\frac{1}{t_{0}}\rfloor-1}$ into a star $K_{1,e(H)+a\lfloor\frac{1}{t_{0}}\rfloor-a}$, where $a=e(H)-|V(H)|+1$.
Through straightforward computation, we readily observe that $\eta(K_{1,e(H)+a\lfloor\frac{1}{t_{0}}\rfloor-a})\leq\eta(H)\leq g_{0}$. Given that there are enough $K_{1,\lfloor\frac{1}{t_{0}}\rfloor-1}$ available, this transformation results in a new graph that satisfies the property $P_{t,r}$. By Lemma 4.3, the star $K_{1,e(H)+a\lfloor\frac{1}{t_{0}}\rfloor-a}$ is walk-preferable to $H\cup aK_{1,\lfloor\frac{1}{t_{0}}\rfloor-1}$. Then, by Lemma 4.1, the transformed graph has a greater spectral radius, leading to a contradiction. Therefore, any non-tree component in $\mathcal{N}_{1}$ must be a $K_{3}$.
By Claim 4.5, $\mathcal{N}_{1}$ cannot contain any non-tree components at all. Consequently, this implies that $\mathcal{N}_{1}$ must be empty. Thus, $G[U]$ is a forest. By Claim 4.2, $G[U]$ consists solely of stars.
Following a similar discussion as in Claim 4.4, by using edge-shifting operations, we can make a star become as large as possible under the constraint of the property $P_{t,r}$. We then classify these stars based on their sizes.
Claim 4.6.
The stars in $G[U]$ can be categorized as follows:
  • Maximal star: $K_{1,\lfloor\frac{1}{t_{0}}\rfloor+a}$, where $a\geq 0$.
  • Big star: $K_{1,\lfloor\frac{1}{t_{0}}\rfloor}$.
  • Small star: $K_{1,\lfloor\frac{1}{t_{0}}\rfloor-1}$.
  • Minimal star: $K_{1,\gamma-1}$, where $1\leq\gamma\leq\lfloor\frac{1}{t_{0}}\rfloor$.
Furthermore, there is at most one maximal (or minimal) star.
These cases are determined by $t_{0}$, $g_{0}$, ensuring $G$ maximizes the spectral radius while satisfying $P_{t,r}$. Notice that when $a=0$, a maximal star reduces to a large star, and when $\gamma=\lfloor\frac{1}{t_{0}}\rfloor-1$, a minimal star transforms into a small star. In particular, when $\gamma=1$, it becomes an isolated vertex.
To maintain consistency with the previous analysis and introduce new notations, we define the following: Let $b$ denote the number of big stars in $G[U]$. Note that when $a=0$, the number of big stars is $b+1$. And let $\phi_{0}=\eta(K_{1,\lfloor\frac{1}{t_{0}}\rfloor})=(1+\lfloor\frac{1}{t_{0}}\rfloor)t_{0}-1$. Clearly, $0<\phi_{0}\leq t_{0}$. We express $g_{0}$ in the form:
$g_{0}=\alpha_{0}\phi_{0}+\beta_{0},$
where $\alpha_{0}$ is a nonnegative integer, and $0\leq\beta_{0}<\phi_{0}$. Let $n-\lfloor t\rfloor-\alpha_{0}=n_{0}\lfloor\frac{1}{t_{0}}\rfloor+\gamma_{0}$, where $1\leq\gamma_{0}\leq\lfloor\frac{1}{t_{0}}\rfloor$. For convenience, we denote $\eta^{+}(G[U])=\sum_{\eta(H)>0}\eta(H)$, where $H$ is a component of $G[U]$. Using these definitions, we now proceed to determine the number of edges in $G[U]$.
@@
Since $G$ satisfies the property $P_{t,r}$, it holds that
$g_{0}\geq\eta^{+}(G[U])=at_{0}+(b+1)\phi_{0}=(a+b+1)\phi_{0}+a(1-t_{0}\lfloor\frac{1}{t_{0}}\rfloor)\geq(a+b+1)\phi_{0}.$
If $a+b+1\geq\alpha_{0}+1$, then we have
$g_{0}\geq\eta^{+}(G[U])\geq(a+b+1)\phi_{0}\geq\alpha_{0}\phi_{0}+\phi_{0}>\alpha_{0}\phi_{0}+\beta_{0}=g_{0},$
which is a contradiction. Therefore, $a+b+1\leq\alpha_{0}$.
Since $G[U]$ is a forest, the number of components is $\big{\lceil}\frac{n-\lfloor t\rfloor-(a+b+1)}{\lfloor\frac{1}{t_{0}}\rfloor}\big{\rceil}$.
Thus, the number of edges in $G[U]$ is
$e(G[U])=n-\lfloor t\rfloor-\big{\lceil}\frac{n-\lfloor t\rfloor-(a+b+1)}{\lfloor\frac{1}{t_{0}}\rfloor}\big{\rceil}.$ (12)
Combining this with $a+b+1\leq\alpha_{0}$, we have
$e(G[U])\leq n-\lfloor t\rfloor-\big{\lceil}\frac{n-\lfloor t\rfloor-\alpha_{0}}{\lfloor\frac{1}{t_{0}}\rfloor}\big{\rceil}=n-\lfloor t\rfloor-(n_{0}+1).$
Next, we show that
$e(G[U])\geq n-\lfloor t\rfloor-\big{\lceil}\frac{n-\lfloor t\rfloor-\alpha_{0}}{\lfloor\frac{1}{t_{0}}\rfloor}\big{\rceil}.$
Suppose, for contradiction, that
$e(G[U])\leq n-\lfloor t\rfloor-\big{\lceil}\frac{n-\lfloor t\rfloor-\alpha_{0}}{\lfloor\frac{1}{t_{0}}\rfloor}\big{\rceil}-1.$
Let $H$ be the union of the maximal star, minimal star, all big stars, and some small stars. By edge-shifting operations, we transform $H$ into a forest $H^{\prime}$, which is the union of $\alpha_{0}$ big stars and a new minimal star. Then, we construct a new graph $G^{\prime}$ with
$\eta^{+}(G^{\prime}[U])=\alpha_{0}\phi_{0}\leq g_{0}.$
This implies that $G^{\prime}$ satisfies the property $P_{t,r}$. Moreover,
$e(G^{\prime}[U])=n-\lfloor t\rfloor-\big{\lceil}\frac{n-\lfloor t\rfloor-\alpha_{0}}{\lfloor\frac{1}{t_{0}}\rfloor}\big{\rceil}\geq e(G[U])+1,$
which implies $\rho(G^{\prime})>\rho(G)$, a contradiction. This completes the proof. ∎
Next, we determine the values of $a$ and $b$. To achieve this, we introduce new parameters. Let $a+b+c+1=\alpha_{0}$, where $c\geq 0$. Since $n-\lfloor t\rfloor=\alpha_{0}+n_{0}\lfloor\frac{1}{t_{0}}\rfloor+\gamma_{0}$, combining with equality (12), we have
$\gamma_{0}+\alpha_{0}-(a+b+1)\leq\lfloor\frac{1}{t_{0}}\rfloor.$
It follows that
$c\leq\lfloor\frac{1}{t_{0}}\rfloor-\gamma_{0}.$
Let $G=G_{(a,b,c)}$, and denote $\eta^{+}(G_{(a,b,c)}[U])=\eta^{+}_{(a,b,c)}$. Clearly,
$\eta^{+}_{(0,\alpha_{0}-1,0)}=\alpha_{0}\phi_{0}=g_{0}-\beta_{0}.$
This implies that $G_{(0,\alpha_{0}-1,0)}$ satisfies the property $P_{t,r}$. Since $G_{(a,b,c)}$ is the spectral extremal graph, it also satisfies the property $P_{t,r}$. Therefore,
$\eta^{+}_{(a,b,c)}=(a+b+1)\phi_{0}+a(t_{0}-\phi_{0})=\alpha_{0}\phi_{0}+at_{0}-(a+c)\phi_{0}\leq g_{0}=\alpha_{0}\phi_{0}+\beta_{0}.$
This yields the following inequalities:
@@
$0\leq c$$\leq$ $\lfloor\frac{1}{t_{0}}\rfloor-\gamma_{0},$ (14)
@@
The inequalities (13), (14), and (15) provide constraints on the parameters $a$, $b$, and $c$, which determine the structure of $G_{(a,b,c)}$. These constraints ensure that $G_{(a,b,c)}$ maximizes the spectral radius while satisfying the property $P_{t,r}$.
Notice that $G_{(a,b,c)}$ can be obtained from $G_{(0,\alpha_{0}-1,0)}$ by the following two types of edge-shifting operations:
  • (I)
    $H_{1}:=(a+1)K_{1,\lfloor\frac{1}{t_{0}}\rfloor}\implies H^{\prime}_{1}:=K_{1,\lfloor\frac{1}{t_{0}}\rfloor+a}\cup aK_{1,\lfloor\frac{1}{t_{0}}\rfloor-1}$.
  • (II)
    $H_{2}:=cK_{1,\lfloor\frac{1}{t_{0}}\rfloor}\cup K_{1,\gamma_{0}-1}\implies H^{\prime}_{2}:=cK_{1,\lfloor\frac{1}{t_{0}}\rfloor-1}\cup K_{1,c+\gamma_{0}-1}$.
We define $f_{k}(a,c)=W_{k}(H_{1}\cup H_{2})-W_{k}(H^{\prime}_{1}\cup H^{\prime}_{2})$. It is easy to see that $f_{1}(a,c)=0$. For $f_{2}(a,c)$, we have
$f_{2}(a,c)$ $=(a+\lfloor\frac{1}{t_{0}}\rfloor)^{2}+(c+\gamma_{0}-1)^{2}+(a+c)(\lfloor\frac{1}{t_{0}}\rfloor-1)^{2}$ (16)
$\quad-(a+c+1)\lfloor\frac{1}{t_{0}}\rfloor^{2}-(\gamma_{0}-1)^{2}$
$=a^{2}+2a\lfloor\frac{1}{t_{0}}\rfloor+c^{2}+2c(\gamma_{0}-1)-2(a+c)\lfloor\frac{1}{t_{0}}\rfloor+(a+c)$
$=a^{2}+a+c^{2}-c-2c\big{(}\lfloor\frac{1}{t_{0}}\rfloor-\gamma_{0}\big{)}.$
We know that the spectral extremal graph is obtained precisely when the function $f(a,c)$ reaches its maximum. Therefore, our goal is to determine the values of $a$ and $c$ that maximize $f(a,c)$.
@@
Let $G_{(a,b,c)}$ be the spectral extremal graph. Then:
$a=\min\big{\{}\big{\lfloor}\frac{\beta_{0}}{t_{0}-\phi_{0}}\big{\rfloor},\alpha_{0}-1\big{\}}\quad\text{and}\quad c=0.$
@@
First, observe that the function $f_{2}(a,c)$ is convex with respect to $a$ and $c$. The symmetry plane of $f(a,c)$ with respect to $c$ is given by:
$c=\big{\lfloor}\frac{1}{t_{0}}\big{\rfloor}-\gamma_{0}+\frac{1}{2}.$
For $0\leq\lfloor\frac{1}{t_{0}}\rfloor-\gamma_{0}<\lfloor\frac{1}{t_{0}}\rfloor-\gamma_{0}+\frac{1}{2}$, the extremum of $f_{2}(a,c)$ must occur at $c=0$.
Next, according to inequalities (13) and (15), we derive the following bounds for $a$ when $c=0$:
$a\leq\frac{\beta_{0}}{t_{0}-\phi_{0}}\quad\text{and}\quad a\leq\alpha_{0}-1.$
Since $a$ is a nonnegative integer and the symmetry plane of $f(a,c)$ with respect to $a$ is $a=-\frac{1}{2}$, it follows that
$a=\min\big{\{}\big{\lfloor}\frac{\beta_{0}}{t_{0}-\phi_{0}}\big{\rfloor},\alpha_{0}-1\big{\}}.$
Note that when $t_{0}=\phi_{0}$, we have $a=\alpha_{0}-1$.
Let $a^{*},b^{*},c^{*}$ be the values of $a,b,c$ that maximize $f_{2}(a,c)$. Then, the optimal values are:
$a^{*}=\min\big{\{}\big{\lfloor}\frac{\beta_{0}}{t_{0}-\phi_{0}}\big{\rfloor},\alpha_{0}-1\big{\}},\quad b^{*}=\alpha_{0}-1-a^{*},\quad c^{*}=0.$
Consequently, the spectral extremal graph $G_{(a^{*},b^{*},c^{*})}$ is obtained. ∎
The spectral extremal graph $G_{(a^{*},b^{*},c^{*})}$ is fully characterized by the optimal values $a^{*}$, $b^{*}$, and $c^{*}$, which are derived from the constraints and the optimization of $f_{2}(a,c)$. This graph maximizes the spectral radius while satisfying the property $P_{t,r}$. By determining the precise structure of $G_{(a^{*},b^{*},c^{*})}$, we complete the structural analysis of the extremal graph $G$.
Recall that $n-\lfloor t\rfloor-\alpha_{0}=n_{0}\lfloor\frac{1}{t_{0}}\rfloor+\gamma_{0}$, where $1\leq\gamma_{0}\leq\lfloor\frac{1}{t_{0}}\rfloor$. Based on the above arguments, when $g_{0}$ does not satisfies that $43t_{0}\leq g_{0}<8t_{0}-$ 4, we have the following theorem.
@@
Let $G$ be the spectral extremal graph that satisfies the property $P_{t,r}$. If $g_{0}$ does not satisfies that $43t_{0}\leq g_{0}<8t_{0}-$ 4, then $G$ is isomorphic to the join of a copy of $K_{\lfloor t\rfloor}$ and a graph $H$, where $H$ consists of stars as follows:
  • One maximal star $K_{1,\lfloor\frac{1}{t_{0}}\rfloor+a}$, where $a=\min\big{\{}\big{\lfloor}\frac{\beta_{0}}{t_{0}-\phi_{0}}\big{\rfloor},\alpha_{0}-1\big{\}}$.
  • $b$ big stars $K_{1,\lfloor\frac{1}{t_{0}}\rfloor}$, where $b=\alpha_{0}-1-a$.
  • $\ell$ small stars $K_{1,\lfloor\frac{1}{t_{0}}\rfloor-1}$, where $\ell=n_{0}-b-1$.
  • One minimal star $K_{1,\gamma-1}$, where $\gamma=\gamma_{0}$.
Note that when $a=0$, it signifies that there is no maximal star in $G[U]$. Consequently, the number of big stars in $G[U]$ simplifies to $\alpha_{0}$. In the case $\gamma_{0}=\lfloor\frac{1}{t_{0}}\rfloor$, it indicates that there is no minimal star present in $G[U]$. The absence of a minimal star changes the composition of small stars. Specifically, the number of small stars becomes $n_{0}-b$.
@@
If $t$ is a positive integer, then $t_{0}=1$. Consequently, we have $\phi_{0}=1$, which implies $a=\alpha_{0}-1$ and $b=0$. Therefore, by applying Theorems 4.1 and 4.2, we obtain the following corollary:
@@
If $t$ is a positive integer, then the spectral extremal graph $G$ is either $G\cong K_{\lfloor t\rfloor}\nabla\big{(}K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}\big{)}$ or $G\cong K_{\lfloor t\rfloor}\nabla\big{(}K_{1,\alpha_{0}}\cup(n-\lfloor t\rfloor-\alpha_{0}-1)K_{1}\big{)}$.

### v3

@@ 4 Characterization of extremal graphs
In the following, we need only consider the case where $g_{0}$ does not satisfy the inequality $43t_{0}\leq g_{0}<8t_{0}-$ 4. By Claims 4.7 and 4.8, $\mathcal{N}_{1}$ must be empty. Thus, $G[U]$ is a forest. Furthermore, $G[U]$ consists solely of stars by Claim 4.4.
According to the definition of $\mathcal{N}_{2}$, the stars within $\mathcal{N}_{2}$ are of the form $K_{1,\lfloor 1/t_{0}\rfloor+a},$ where $a\geq 0$. By Claim 4.6, there are only two distinct types of stars in $\mathcal{N}_{2}$. We refer to the star as a big star when $a=0$, and as a maximal star for cases where $a\geq 1$. For $\mathcal{N}_{3}$, the stars are of the form $K_{1,\gamma-1}$, where $1\leq\gamma\leq\lfloor 1/t_{0}\rfloor$. These stars are referred to as small stars when $\gamma=\lfloor 1/t_{0}\rfloor$, and as minimal stars when $1\leq\gamma<\lfloor 1/t_{0}\rfloor$. By Claims 4.5 and 4.6, there exists at most one minimal (or maximal) star. In particular, when $\gamma=1$, a minimal star becomes an isolated vertex.
To determine the number of edges in $G[U]$, we introduce the following notations: Let $b$ denote the number of big stars (i.e., $K_{1,\lfloor 1/t_{0}\rfloor}$) in $G[U]$. For convenience, when considering the maximal star $K_{1,\lfloor 1/t_{0}\rfloor+a}$, allow $a=0$. If $a=0$, then the number of big stars is $b+1$; otherwise, there is exactly one maximal star of the form $K_{1,\lfloor 1/t_{0}\rfloor+a}$ with $a\geq 1$.
Define $\phi_{0}=\eta(K_{1,\lfloor 1/t_{0}\rfloor})=(1+\lfloor 1/t_{0}\rfloor)t_{0}-1$, noting that $0<\phi_{0}\leq t_{0}$. Express $g_{0}$ as
$g_{0}=\alpha_{0}\phi_{0}+\beta_{0},$ (19)
where $\alpha_{0}$ is a nonnegative integer and $0\leq\beta_{0}<\phi_{0}$.
Additionally, let $n-\lfloor t\rfloor-\alpha_{0}=n_{0}\lfloor 1/t_{0}\rfloor+\gamma_{0}$, where $n_{0}$ is an integer sufficiently large, and $1\leq\gamma_{0}\leq\lfloor 1/t_{0}\rfloor$. Using these definitions, we now proceed to determine $e(G[U])$.
@@
By Claim 4.3, the stars within $\mathcal{N}_{2}$ are positive. Moreover, recall that $\mathcal{N}_{3}$ contains no positive stars. Then, $\eta^{+}(G[U])=\eta(K_{1,\lfloor 1/t_{0}\rfloor+a})+b\eta(K_{1,\lfloor 1/t_{0}\rfloor}).$ It follows that
$\eta^{+}(G[U])=at_{0}+(b+1)\phi_{0}=(a+b+1)\phi_{0}+a(t_{0}-\phi_{0})\geq(a+b+1)\phi_{0}.$ (20)
Note that $\eta^{+}(G[U])\leq g_{0}$. If $a+b+1\geq\alpha_{0}+1$, then $g_{0}\geq(\alpha_{0}+1)\phi_{0}>\alpha_{0}\phi_{0}+\beta_{0}=g_{0}$, which is a contradiction. Therefore,
$a+b+1\leq\alpha_{0}.$ (21)
Since $G[U]$ is a forest, and based on the classification of $\mathcal{N}_{2}\cup\mathcal{N}_{3}$, the number of components in $G[U]$ is $\big{\lceil}\frac{n-\lfloor t\rfloor-(a+b+1)}{\lfloor 1/t_{0}\rfloor}\big{\rceil}$. Thus, the number of edges in $G[U]$ is
$e(G[U])=n-\lfloor t\rfloor-\Big{\lceil}\frac{n-\lfloor t\rfloor-(a+b+1)}{\lfloor 1/t_{0}\rfloor}\Big{\rceil}.$ (22)
Combining (21) and (22), we obtain that
$e(G[U])\leq n-\lfloor t\rfloor-\Big{\lceil}\frac{n-\lfloor t\rfloor-\alpha_{0}}{\lfloor 1/t_{0}\rfloor}\Big{\rceil}=n-\lfloor t\rfloor-(n_{0}+1).$
Next, we show that $e(G[U])\geq n-\lfloor t\rfloor-(n_{0}+1)$. Suppose, to the contrary, that this is not the case. Let $F$ be the union of the maximal star, the minimal star, all big stars, and some (finite) small stars. By edge-shifting operations, we can transform $F$ into a forest $F^{\prime}$, which is the union of $\alpha_{0}$ big stars and a new minimal star. Then construct a graph $G^{\prime}$ from $G$ by replacing the edges of $F$ with the edges of $F^{\prime}$. It is clear that $\eta^{+}(G^{\prime}[U])=\alpha_{0}\phi_{0}\leq g_{0}$, which implies that $G^{\prime}$ satisfies the property $P_{t,r}$. Moreover,
$e(G^{\prime}[U])=n-\lfloor t\rfloor-\Big{\lceil}\frac{n-\lfloor t\rfloor-\alpha_{0}}{\lfloor 1/t_{0}\rfloor}\Big{\rceil}=n-\lfloor t\rfloor-(n_{0}+1)\geq e(G[U])+1.$
By Corollary 4.1, we have $\rho(G^{\prime})>\rho(G)$, a contradiction. This completes the proof. ∎
Next, we determine the values of $a$ and $b$. To achieve this, we shall introduce a new parameter. Denote $c=\alpha_{0}-(a+b+1)$. Based on (21), it follows that $c\geq 0$. Recall that $n-\lfloor t\rfloor=\alpha_{0}+\gamma_{0}+n_{0}\lfloor 1/t_{0}\rfloor$. Combining this with equality (22) and Claim 4.9, we have $\alpha_{0}+\gamma_{0}-(a+b+1)\leq\lfloor 1/t_{0}\rfloor$. It follows that $c\leq\lfloor 1/t_{0}\rfloor-\gamma_{0}$.
We can observe that $G$ is uniquely determined by $a$ and $b$. Therefore, we may denote $G=G_{(a,b,c)}$, and let $\eta^{+}(G_{(a,b,c)}[U])=\eta^{+}_{(a,b,c)}$. Recall that $\eta^{+}_{(a,b,c)}=\eta^{+}(G[U])\leq g_{0}$. Furthermore, based on (20), we obtain that
$\eta^{+}_{(a,b,c)}=at_{0}+(b+1)\phi_{0}=at_{0}+(\alpha_{0}-a-c)\phi_{0}\leq g_{0}=\alpha_{0}\phi_{0}+\beta_{0}.$ (23)
From the previous discuss, we get the following inequalities:
@@
$0\leq c$$\leq$ $\lfloor 1/t_{0}\rfloor-\gamma_{0},$ (25)
@@
In view of (23), we have $\eta^{+}_{(0,\alpha_{0}-1,0)}=\alpha_{0}\phi_{0}=g_{0}-\beta_{0},$ and by Lemma 4.5, we see that $G_{(0,\alpha_{0}-1,0)}$ satisfies the property $P_{t,r}$. Observe that $G_{(0,\alpha_{0}-1,0)}$ contains $\alpha_{0}$ components of the form $K_{1,\lfloor 1/t_{0}\rfloor}$. Additionally, recall that $n-\lfloor t\rfloor-n_{0}\lfloor 1/t_{0}\rfloor=\alpha_{0}+\gamma_{0}$. This further implies that $G_{(0,\alpha_{0}-1,0)}$ contains a minimal component $K_{1,\gamma_{0}-1}$ and $n_{0}-\alpha_{0}$ components of the form $K_{1,\lfloor 1/t_{0}\rfloor-1}$. Recall that $\alpha_{0}=a+b+c+1$. It is easy to see that $G_{(a,b,c)}$ can be obtained from $G_{(0,\alpha_{0}-1,0)}$ by the following two types of edge-shifting operations:
(i) $F_{1}:=(a+1)K_{1,\lfloor 1/t_{0}\rfloor}\implies F^{\prime}_{1}:=K_{1,\lfloor 1/t_{0}\rfloor+a}\cup aK_{1,\lfloor 1/t_{0}\rfloor-1}$;
(ii) $F_{2}:=cK_{1,\lfloor 1/t_{0}\rfloor}\cup K_{1,\gamma_{0}-1}\implies F^{\prime}_{2}:=cK_{1,\lfloor 1/t_{0}\rfloor-1}\cup K_{1,c+\gamma_{0}-1}$.
We now define $f_{k}(a,c)=W_{k}(F^{\prime}_{1}\cup F^{\prime}_{2})-W_{k}(F_{1}\cup F_{2})$. Then, it is clear that $f_{1}(a,c)=0$ since $e(F^{\prime}_{1}\cup F^{\prime}_{2})=e(F_{1}\cup F_{2})$. Note that $W_{2}(F)=\sum_{v\in V(F)}\big{(}d_{F}(v)\big{)}^{2}$. We further have
$f_{2}(a,c)$ $=(a+\lfloor 1/t_{0}\rfloor)^{2}+(c+\gamma_{0}-1)^{2}+(a+c)(\lfloor 1/t_{0}\rfloor-1)^{2}$
$\quad-(a+c+1)\lfloor 1/t_{0}\rfloor^{2}-(\gamma_{0}-1)^{2}$
$=a^{2}+2a\lfloor 1/t_{0}\rfloor+c^{2}+2c(\gamma_{0}-1)-2(a+c)\lfloor 1/t_{0}\rfloor+(a+c)$
$=a^{2}+a+c^{2}-2c\big{(}\lfloor 1/t_{0}\rfloor-\gamma_{0}+1/2\big{)}.$
Subject to the property $P_{t,r}$, the extremal graph $G$ is obtained precisely when the function $f_{2}(a,c)$ reaches its maximum. Therefore, our goal is to determine the values of $a$ and $c$ that maximize $f_{2}(a,c)$, under the condition that $\eta^{+}_{(a,b,c)}\leq g_{0}$.
@@
$a=\min\big{\{}\big{\lfloor}\frac{\beta_{0}}{t_{0}-\phi_{0}}\big{\rfloor},\alpha_{0}-1\big{\}}$, $b=\alpha_{0}-a-1$ and $c=0$.
@@
First, observe that $f_{2}(a,c)$ is a quadratic function with respect to $a$ and $c$, and it represents an upward-opening paraboloid of revolution. The symmetry plane of $f(a,c)$ with respect to $c$ is $c=\big{\lfloor}1/t_{0}\big{\rfloor}-\gamma_{0}+\frac{1}{2}$. Form (25), we know that $0\leq c\leq\lfloor 1/t_{0}\rfloor-\gamma_{0}$. Then, the extremum of $f_{2}(a,c)$ must occur when $c=0$.
Next, from (24) and (26), we can derive that $a\leq\frac{\beta_{0}}{t_{0}-\phi_{0}}$ and $a\leq\alpha_{0}-1$ when $c=0$. Since $a$ is a nonnegative integer and the symmetry plane of $f_{2}(a,c)$ with respect to $a$ is $a=-\frac{1}{2}$, it follows that $a=\min\big{\{}\big{\lfloor}\frac{\beta_{0}}{t_{0}-\phi_{0}}\big{\rfloor},\alpha_{0}-1\big{\}},$ where $a=\alpha_{0}-1$ when $t_{0}=\phi_{0}$. Recall that $c=\alpha_{0}-(a+b+1)$. We further have $b=\alpha_{0}-a-1$.
Now it suffices to verify that $\eta^{+}_{(a,b,c)}\leq g_{0}$ for the values of $a,b,c$ obtained above. In view of (23), we have
$\eta^{+}_{(a,b,c)}=at_{0}+(b+1)\phi_{0}=\alpha_{0}\phi_{0}+a(t_{0}-\phi_{0})\leq\alpha_{0}\phi_{0}+\beta_{0}=g_{0}.$
This completes the proof. ∎
Recall that $n-\lfloor t\rfloor-\alpha_{0}=n_{0}\lfloor 1/t_{0}\rfloor+\gamma_{0}$, where $1\leq\gamma_{0}\leq\lfloor 1/t_{0}\rfloor$. Based on the above arguments, the following theorem holds.
@@
Let $G$ be a graph that attains the maximum spectral radius over all $n$-vertex graphs satisfying the property $P_{t,r}$. If $g_{0}$ does not satisfies $43t_{0}\leq g_{0}<8t_{0}-$ 4, then $G\cong K_{\lfloor t\rfloor}\nabla H$, where $H$ consists of stars as follows:
(i) one maximal star $K_{1,\lfloor 1/t_{0}\rfloor+a}$, where $a=\min\big{\{}\big{\lfloor}\frac{\beta_{0}}{t_{0}-\phi_{0}}\big{\rfloor},\alpha_{0}-1\big{\}}$;
(ii) $b$ big stars $K_{1,\lfloor 1/t_{0}\rfloor}$, where $b=\alpha_{0}-a-1$;
(iii) $\ell$ small stars $K_{1,\lfloor 1/t_{0}\rfloor-1}$, where $\ell=n_{0}-b-1$;
(iv) one minimal star $K_{1,\gamma-1}$, where $\gamma=\gamma_{0}$.
Note that when $a=0$, it indicates that there is no maximal star in $G[U]$. Consequently, the number of big stars in $G[U]$ simplifies to $\alpha_{0}$. In the case $\gamma_{0}=\lfloor 1/t_{0}\rfloor$, it implies that there is no minimal star in $G[U]$. The absence of a minimal star changes the composition of small stars. Specifically, the number of small stars becomes $n_{0}-b$.
@@
If $t$ is a positive integer, then $t_{0}=\lfloor t\rfloor+1-t=1$ and $\phi_{0}=(1+\lfloor 1/t_{0}\rfloor)t_{0}-1=1$. Based on (19), we further have $\alpha_{0}=\lfloor g_{0}\rfloor.$ It follows that $a=\alpha_{0}-1=\lfloor g_{0}\rfloor-1$ and $b=0$. By applying Theorems 4.1 and 4.2, we obtain the following corollary.
@@
Let $G$ be a graph that attains the maximum spectral radius over all $n$-vertex graphs satisfying the property $P_{t,r}$. If $t$ is a positive integer, then $G\cong K_{t}\nabla\big{(}K_{3}\cup(n-t-3)K_{1}\big{)}$ for $43\leq g_{0}<$ 4, or $G\cong K_{t}\nabla\big{(}K_{1,\lfloor g_{0}\rfloor}\cup(n-t-\lfloor g_{0}\rfloor-1)K_{1}\big{)}$ otherwise.
For $t$ being a positive integer, Guiduli proved the Theorem 1.1, and he also conjectured that the spectral extremal graph with respect to Theorem 1.1 is $K_{t}\nabla F$ where the edges within $F$ form a star. Corollary 4.3 resolves this conjecture, and provides a complete characterization of the extremal graphs.